OpenStudy (anonymous):

Integrate sin^9(x)dx

OpenStudy (anonymous):

\[\int\limits_{}^{}sinx.dx = -cosx + c\]

OpenStudy (anonymous):

@waterineyes i think u made a little mistake

OpenStudy (anonymous):

Correct me dear...

OpenStudy (unklerhaukus):

\[\int\sin^nx\cdot\text dx=-\frac1n\cos\sin^{n-1}x+\frac{n-1}{n}\int\sin^{n-2}x\cdot\text dx\]

OpenStudy (anonymous):

Oh sorry then my mistake...

OpenStudy (anonymous):